# Deep in the Knotted Black Hole

### v2

@@ Abstract
We consider the transition rate of a freely falling Unruh-DeWitt (UDW) detector, coupled linearly to a massless scalar quantum field prepared in the Hartle-Hawking-Israel state, as a probe of the interior of a black hole. Specifically, we consider the transition rate of a detector in the spinless Bañados-Teitelboim-Zanelli (BTZ) black hole as it freely falls toward and across the horizon and compare it to the corresponding situation for an $\mathbb{R}\text{P}^{2}$ geon. Both the BTZ black hole and its geon counterpart are quotients of $\text{AdS}_{3}$ spacetime that are identical exterior to the horizon but have different interior topologies. We find outside the horizon that the rates are qualitatively similar, but with the amplitude in the geon spacetime larger than in the BTZ case. Once the detector crosses the horizon, there are notable distinctions characterized by different discontinuities in the temporal derivative of the response rate. These discontinuities can appear outside the horizon if the detector is switched on at a sufficiently early time, within the past white hole horizon. In general, the detector can act as an ‘early warning system’ that both spots the black hole horizon and discerns its interior topology.
@@ I Introduction
Studies in $2+1$ dimensions have the advantage that one can work with exact solutions to the Einstein equations, unlike the aforementioned $(1+1)$-dimensional settings; furthermore, the detector can be coupled linearly to the field, instead of its momentum, without infrared complications. In spacetimes that are quotients of a maximally symmetric space, there is a further technical advantage insofar as the Wightman function of the quantum scalar field can be computed as a sum over images instead of the unavoidable mode sum in the Schwarzschild case [22]. It was shown some time ago [9] that a UDW detector radially falling toward a static BTZ black hole exhibits the expected smooth and non-thermal response when the detector is still outside the horizon. Recently, this study was extended to situations where the detector crosses the horizon [28]. The response rate of the detector was found to be a smooth and slowly oscillating function, punctuated by non-differentiable points denoted as ‘glitches’. Similar behaviour was found for the response rate of a detector falling into a rotating BTZ black hole [29]. In this case four types of glitches were observed, with their locations dependent on the mass and angular momentum of the black hole. Taken together, these results suggest that the event horizon of a black hole may be discernable to a local probe when quantum-field theoretic effects are included.
Here we investigate the effects of topology hidden behind the horizon on the transition rate of a freely falling detector before, at, and after the detector crosses the horizon. We consider the topological geon black hole spacetime known as the $\mathbb{R}\text{P}^{2}$ geon [11], which is a $\mathbb{Z}_{2}$ quotient of the spinless $(2+1)$-dimensional Bañados-Teitelboim-Zanelli (BTZ) black hole, to which it is locally isometric, as well as being locally isometric to $(2+1)$-dimensional anti-de Sitter (AdS) spacetime.
@@ III.1 BTZ glitches
The argument of the square root in the $\zeta$-dependent term is positive for all $\tilde{s}$ in the interval $20<\tilde{s}\leq\Delta\tilde{\tau}<\pi/$ 2 and all $n\geq 1$. However in the first term, the argument of the square root can change of sign within the interval of integration for $n\geq 1$. This results in a singularity in the integrand, but the singularity is of the inverse square root type and hence integrable. Analytically solving the equation $\sigma^{\text{BTZ}}_{n}(\tau,\tau-s)=0$, we obtain
@@ III.2 Geon glitches
The transition rate associated with the geon black hole consists of two terms. One term is identical to the BTZ rate (23); the other is given by (25). This extra term is obtained from the BTZ rate (23) by substitution of (20b) in place of (20a). This means that the glitches present in the transition rate of a detector in the geon spacetime will be the same as those in the BTZ case, plus additional glitches specific to the geon, obtained by solving $\sigma^{\text{geon}}_{n}(\tau^{geon}_{n},\tau_{0})=0$, or, for $\zeta=\pm 1$, also $\sigma^{\text{geon}}_{n}(\tau^{geon}_{n},\tau_{0})+2=0$. Geometrically, the second case corresponds to a null geodesic that connects the switch-on event to a later event on the trajectory after reflection from infinity. In this section we consider only the first of these cases, but we shall return to the second case in Section IV.4. We note from (22b) that the sum in (25) is invariant under $n\to-1-n$, so it suffices to sum over $n\geq 0$ and double the result.
@@ IV.2 $t_{0}\neq 0=\tau_{0}$
Figure 14: Transition rate in the geon spacetimes for $\tau_{0}=0$, $t_{0}/\ell=-200$, $q=1.2$, $M=10^{-4}$, $\Omega\ell=-5$ and $\zeta\in\{1,0,-1\}$. The vertical dashed line is the horizon, located at $\tau/\ell=\arccos(1/q)=0.5856$. The black dots are the glitches from the $\zeta$-independent term, at (31). The red dots are the glitches from the term proportional to $\zeta$, at (38); as discussed around (39), these are present only for $t_{0}/\ell<t_{0}^{b,{lim}}/\ell\approx-59.947$.
@@ IV.3 $t_{0}=0\neq\tau_{0}$
The figures show that glitches move towards earlier proper times as $\tau_{0}$ becomes more negative. Glitches before crossing the black hole horizon can occur only when $\tau_{0}<-\tau_{H}$, so that the detector has started to operate before emerging from the white hole part of the spacetime, as shown in Figure 4 lower panel: one such glitch is seen for one of the trajectories in Figure 11.
@@ IV.4 $\zeta\neq 0$
However, there are regions of the parameter space where the $\zeta=\pm 1$ term introduces new glitches in the response. Geometrically, these new glitches come from null geodesics that connect the switch-on moment to a later moment on the trajectory by reflection from infinity. Consideration of Figure 3 lower panel shows that one situation where this happens is for $t_{0}<0=\tau_{0}$ when $t_{0}$ is sufficiently negative, and the new glitches then occur after the detector has entered the black hole. Analytically, the new glitches come by solving $\sigma_{n}^{\text{geon}}\left(\tau^{b}_{n},0\right)+2=0$, which yields a new set of glitches positioned at
@@
where the superscript ‘b’ denotes that these glitches originate from the $\zeta$ term in (25). These glitches exist for $t_{0}<t_{0}^{b,{lim}}$, where
$t_{0}^{b,{lim}}=-\frac{\ell}{2\sqrt{M}}\operatorname{arccosh}\!\left({\frac{q}{\sqrt{q^{2}-1}}}\right).$ (39)

### v3

@@ Abstract
We consider the transition rate of a freely falling Unruh-DeWitt detector, coupled linearly to a massless scalar quantum field prepared in the Hartle-Hawking-Israel state, as a probe of the interior of a black hole. Specifically, we consider the transition rate of a detector in the spinless Bañados-Teitelboim-Zanelli (BTZ) black hole as it freely falls toward and across the horizon and compare it to the corresponding situation for an $\mathbb{R}\text{P}^{2}$ geon. Both the BTZ black hole and its geon counterpart are quotients of $\text{AdS}_{3}$ spacetime that are identical exterior to the horizon but have different interior topologies. We find outside the horizon that the rates are qualitatively similar, but with the amplitude in the geon spacetime larger than in the BTZ case. Once the detector crosses the horizon, there are notable distinctions characterized by different discontinuities in the temporal derivative of the response rate. These discontinuities can appear outside the horizon if the detector is switched on at a sufficiently early time, within the past white hole horizon. In general, the detector can act as an ‘early warning system’ that both spots the black hole horizon and discerns its interior topology.
@@ I Introduction
Studies in $2+1$ dimensions have the advantage that one can work with exact solutions to the Einstein equations, unlike the aforementioned $(1+1)$-dimensional settings; furthermore, the detector can be coupled linearly to the field, instead of its momentum, without infrared complications. In spacetimes that are quotients of a maximally symmetric space, there is a further technical advantage insofar as the Wightman function of the quantum scalar field can be computed as a sum over images instead of the unavoidable mode sum in the Schwarzschild case [22, 23]. It was shown some time ago [9] that a UDW detector radially falling toward a static Bañados-Teitelboim-Zanelli (BTZ) black hole exhibits the expected smooth and non-thermal response when the detector is still outside the horizon. Recently, this study was extended to situations where the detector crosses the horizon [28]. The response rate of the detector was found to be a smooth and slowly oscillating function, punctuated by non-differentiable points denoted as ‘glitches’. Similar behaviour was found for the response rate of a detector falling into a rotating BTZ black hole [29]. In this case four types of glitches were observed, with their locations dependent on the mass and angular momentum of the black hole. Taken together, these results suggest that the event horizon of a black hole may be discernable to a local probe when quantum-field theoretic effects are included.
Here we investigate the effects of topology hidden behind the horizon on the transition rate of a freely falling detector before, at, and after the detector crosses the horizon. We consider the topological geon black hole spacetime known as the $\mathbb{R}\text{P}^{2}$ geon [11], which is a $\mathbb{Z}_{2}$ quotient of the spinless $(2+1)$-dimensional BTZ black hole, to which it is locally isometric, as well as being locally isometric to $(2+1)$-dimensional anti-de Sitter (AdS) spacetime.
@@ III.1 BTZ glitches
The argument of the square root in the $\zeta$-dependent term is positive for all $\tilde{s}$ in the interval $20<\tilde{s}\leq\Delta\tilde{\tau}<\pi/$ 2 and all $n\geq 1$. However in the first term, the argument of the square root can change sign within the interval of integration for $n\geq 1$. This results in a singularity in the integrand, but the singularity is of the inverse square root type and hence integrable. Analytically solving the equation $\sigma^{\text{BTZ}}_{n}(\tau,\tau-s)=0$, we obtain
@@ III.2 Geon glitches
The transition rate associated with the geon black hole consists of two terms. One term is identical to the BTZ rate (23); the other is given by (25). This extra term is obtained from the BTZ rate (23) by substitution of (20b) in place of (20a). This means that the glitches present in the transition rate of a detector in the geon spacetime will be the same as those in the BTZ case, plus additional glitches specific to the geon, obtained by solving $\sigma^{\text{geon}}_{n}(\tau^{\text{geon}}_{n},\tau_{0})=0$, or, for $\zeta=\pm 1$, also $\sigma^{\text{geon}}_{n}(\tau^{\text{geon}}_{n},\tau_{0})+2=0$. Geometrically, the second case corresponds to a null geodesic that connects the switch-on event to a later event on the trajectory after reflection from infinity. In this section we consider only the first of these cases, but we shall return to the second case in Section IV.4. We note from (22b) that the sum in (25) is invariant under $n\to-1-n$, so it suffices to sum over $n\geq 0$ and double the result.
@@ IV.2 $t_{0}\neq 0=\tau_{0}$
Figure 14: Transition rate in the geon spacetimes for $\tau_{0}=0$, $t_{0}/\ell=-200$, $q=1.2$, $M=10^{-4}$, $\Omega\ell=-5$ and $\zeta\in\{1,0,-1\}$. The vertical dashed line is the horizon, located at $\tau/\ell=\arccos(1/q)=0.5856$. The black dots are the glitches from the $\zeta$-independent term, at (31). The red dots are the glitches from the term proportional to $\zeta$, at (38); as discussed around (39), these are present only for $t_{0}/\ell<t_{0}^{\text{b},\text{lim}}/\ell\approx-59.947$.
@@ IV.3 $t_{0}=0\neq\tau_{0}$
The figures show that glitches move toward earlier proper times as $\tau_{0}$ becomes more negative. Glitches before crossing the black hole horizon can occur only when $\tau_{0}<-\tau_{H}$, so that the detector has started to operate before emerging from the white hole part of the spacetime, as shown in Figure 4 lower panel: one such glitch is seen for one of the trajectories in Figure 11.
@@ IV.4 $\zeta\neq 0$
However, there are regions of the parameter space where the $\zeta=\pm 1$ term introduces new glitches in the response. Geometrically, these new glitches come from null geodesics that connect the switch-on moment to a later moment on the trajectory by reflection from infinity. Consideration of Figure 3 lower panel shows that one situation where this happens is for $t_{0}<0=\tau_{0}$ when $t_{0}$ is sufficiently negative, and the new glitches then occur after the detector has entered the black hole. Analytically, the new glitches come by solving $\sigma_{n}^{\text{geon}}\left(\tau^{\text{b}}_{n},0\right)+2=0$, which yields a new set of glitches positioned at
@@
where the superscript ‘b’ denotes that these glitches originate from the $\zeta$ term in (25). These glitches exist for $t_{0}<t_{0}^{\text{b},\text{lim}}$, where
$t_{0}^{\text{b},\text{lim}}=-\frac{\ell}{2\sqrt{M}}\operatorname{arccosh}\!\left({\frac{q}{\sqrt{q^{2}-1}}}\right).$ (39)